OpenStudy (lena772):

Using your function f(x) from question 2. Demonstrate and explain how to find the average rate of change between year 3 and 5, and between year 5 and 7. Explain what the average rate of change represents to the frog population.

OpenStudy (lena772):

\[f(x) = A * (B)^x\]

OpenStudy (lena772):

@ranga

OpenStudy (lena772):

@BigTeach2345 Think you can help?

OpenStudy (lena772):

@AllTehMaffs @ganeshie8 @shamil98 @hartnn

ganeshie8 (ganeshie8):

assign some values to A and B

ganeshie8 (ganeshie8):

initial frog population = A = 10 ? rate of change of population = B = 2 times each year ?

OpenStudy (lena772):

A=10 B=2

OpenStudy (lena772):

f(x)=10*2^x

ganeshie8 (ganeshie8):

looks good ! ``` Demonstrate and explain how to find the average rate of change between year 3 and 5, ```

ganeshie8 (ganeshie8):

wat do u knw about average rate of change ?

OpenStudy (lena772):

the slope stays the same right?

ganeshie8 (ganeshie8):

yes average rate of change of \(f(x)\) between \(x = a\) and \(x = b\) is :- \(\huge \frac{f(b) -f(a)}{b-a}\)

ganeshie8 (ganeshie8):

thats the formula, see if u can apply it, and interpret wat the formula means.

OpenStudy (lena772):

f(2)-f(10)/2-10

OpenStudy (lena772):

f(2)=10*2^2 f(2)=10*4 f(2)= 40?

OpenStudy (lena772):

f(10)=10*2^10 f(10)=10*1024 f(10=10240?

ganeshie8 (ganeshie8):

yup ! keep going

OpenStudy (lena772):

40-10240/2-10

OpenStudy (lena772):

-10200/-8

OpenStudy (lena772):

1275?

ganeshie8 (ganeshie8):

yes, wat units ?

ganeshie8 (ganeshie8):

1275 frogs per .... ?

OpenStudy (lena772):

1275 frogs per year?

ganeshie8 (ganeshie8):

in ur previous question f(x) = A*B^x, x is in years is it ?

OpenStudy (lena772):

yes

ganeshie8 (ganeshie8):

then yes, its 1275 frogs per year.

ganeshie8 (ganeshie8):

now explain in ur own words, wat the average rate 1275 represents. use 2-3 sentences

OpenStudy (lena772):

But I am supposed to be looking between the years 3 and 5 and years 5 and 7...

ganeshie8 (ganeshie8):

yes it is defenitely an increase, but why did u calculate the average rate between x = 2 and x = 10 ? the question is asking u to calculate between x=3 and x=5 right ?

ganeshie8 (ganeshie8):

yes.... lol why did we do average rate between x = 2 and x = 10 :o

OpenStudy (lena772):

idk :'(

OpenStudy (lena772):

And also x=5 and x=7

ganeshie8 (ganeshie8):

its okay, compute the average rate again between 3 & 5 and, 5 and 7 more practice ! good for u :D

OpenStudy (lena772):

f(b)-f(a)/b-a f(3)-f(5)/3-5 f(3)=?

OpenStudy (lena772):

Are we still using 10 and 2?

ganeshie8 (ganeshie8):

we're done wid 10 and 2 right. now you're doing average rate between 3 & 5

OpenStudy (lena772):

What are A and B when x=3 ?

ganeshie8 (ganeshie8):

\(f(x) = 10*2^x\)

OpenStudy (lena772):

so f(3)=10*2^3 f(3)= 10*8 f(3) = 80

ganeshie8 (ganeshie8):

Yes !

OpenStudy (lena772):

f(5)=10*2^5 f(5)=10*32 f(5)=320

ganeshie8 (ganeshie8):

yes ! keep going..

OpenStudy (lena772):

80-320/2-10 240/-2 -120

OpenStudy (lena772):

I'm confused :/

ganeshie8 (ganeshie8):

80-320/2-10 -240/-2 120

OpenStudy (lena772):

Right, okay. But what I don't understand is the equation is f(b)-f(a)/b-a? how can f(b)=3 and f(a)=5 but a=10 and b =2?

ganeshie8 (ganeshie8):

forget about A and B here. your function is simply : \(f(x) = 10*2^x\)

OpenStudy (lena772):

Ok

ganeshie8 (ganeshie8):

average rate of change between \(x = m\) and \(x = n \) is :- \(\huge \frac{f(m) - f(n)}{m-n}\)

ganeshie8 (ganeshie8):

if it helps, try to use above formula. i thinkg a and b are confusing us

OpenStudy (lena772):

what are m and n

ganeshie8 (ganeshie8):

\(f(x) = 10*2^x\) average rate of change between \(x = 3\) and \(x = 5\) is :- \(\huge \frac{f(3) - f(5)}{3-5}\)

OpenStudy (lena772):

-240/-2 120

OpenStudy (lena772):

120 frogs per year

ganeshie8 (ganeshie8):

yes !

OpenStudy (lena772):

Ok so now 5 and 7

OpenStudy (lena772):

f(5)-f(7)/5-7 f(5)-f(7)/-2

OpenStudy (lena772):

f(5)=10*2^5 f(5)=10*32 f(5) = 320 f(7)=10*2^7 f(7)=10*128 f(7)=1280

OpenStudy (lena772):

320-1280/-2 -960/-2 480

OpenStudy (lena772):

480 frogs per year?

OpenStudy (lena772):

How come it is different? :S

ganeshie8 (ganeshie8):

good question, may be u can try adn tell me why the change is not same :)

ganeshie8 (ganeshie8):

hint : our function is not a straight line, to have FIXED slope

OpenStudy (lena772):

Oh because as the time increases so does the populace, directly proportional?

ganeshie8 (ganeshie8):

nope. u doing calculus right. use that knowledge

ganeshie8 (ganeshie8):